OpenStudy (anonymous):

how to fully factorize: 5(x+2) -x(x+2)

OpenStudy (nirmalnema):

1st open the bracket

OpenStudy (nirmalnema):

u will get.....5x+10-x^2-2x = multiply it with - sign so u get x^2-3x-10 use spliting the middle term x^2-5x+2x-10 x(x-5)+2(x-5) (x-5)(x+2) x=5 and -2

OpenStudy (anonymous):

There is a common factor that is (x+2). You can do 5(x+2)-x(x+2)=(x+2)(5-x)